OpenStudy (anonymous):

I'LL FAN AND GIVE YOU A MEDAL IF ANSWER THIS QUESTIONS 2. Find the volume of a sphere with a radius of 10 feet. Round to the nearest whole number. 1,257 ft3 4,187 ft3 1,089 ft3 1,568 ft3 3. Find the surface area for a sphere with a radius of 7 cm. Round to the nearest whole number. 307 cm2 1,436 cm2 1,020 cm2 615 cm2 4. Find the volume for a sphere with a radius of 7 cm. Round to the nearest whole number. 307 cm2 1,436 cm2 1,020 cm2 615 cm2 5. Find the radius of a sphere with a surface area of 804 cm2. (1 point) 9 cm 8 cm 64 cm 204 cm

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@Lcb2516

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Find the volume of a sphere with a radius of 10 feet. Round to the nearest whole number. 1,257 ft3 4,187 ft3 1,089 ft3 1,568 ft3 Okay, so the formula for the volume of a sphere is: \(V = \dfrac{4}{3} \pi r^3\) Plug in what we know: \(V = \dfrac{4}{3} (3.14)(10^3)\) Can you multiply 4/3 * 3.14?

OpenStudy (anonymous):

Real quick, is that a test and youare looking to cheat? If you're not cheating I could possibly answer some of these

OpenStudy (anonymous):

@Vanquish If by 'answer' you mean 'give him the answer' we are not here to do that..we have to guide them to the answer, and we need their participation.

OpenStudy (anonymous):

@Vanquish is a quick chek

OpenStudy (anonymous):

@iGreen. Alright, I am new to this so please forgive me if I do not do things corectly

OpenStudy (anonymous):

its ok :) @Vanquish

OpenStudy (anonymous):

Yes, it's perfectly fine as you are new. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Now, can you multiply 4/3 * 3.14? @bmo1345

OpenStudy (anonymous):

how do u * a fraction on the calculator? @iGreen.

OpenStudy (anonymous):

Google works, but First you divide 4 / 3

OpenStudy (anonymous):

Then you multiply 3.14

OpenStudy (anonymous):

4.18....

OpenStudy (anonymous):

Yes, now multiply by 10^3, which is 1000 What's 1000 * 4.18?

OpenStudy (anonymous):

10^3 is 30

OpenStudy (anonymous):

No, it's 1000. \(10^3 \rightarrow 10 \times 10 \times 10\)

OpenStudy (anonymous):

but 1000* 4.18= 4180.0

OpenStudy (anonymous):

Yes, but when you multiplied 4/3 * 3.14, the actual result is: 4.18666667 And 4.18666667 * 1000 = 4186.666667 And it says to round to the nearest whole number so we have: 4187

OpenStudy (anonymous):

so #2 is B

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\sf\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!!!}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

Btw, we are only allowed to ask 1 question per post, so close this post and open a new one.

OpenStudy (anonymous):

I have to go now..so someone else will view your question and help.

OpenStudy (anonymous):

@Vanquish can u give me the quick answer to them plz and i'll fan you

OpenStudy (anonymous):

iGreen said I can't give an answer but I can guild you

OpenStudy (anonymous):

com' n plz